OpenStudy (anonymous):

Washburn is seeking a Sales and Marketing Coordinator with a bachelor's degree or equivalent experience,knowledge in Microsoft Office. This position poays $25,000 to $35,000, depending on experience. Assume a person is paid weekly and earns $32,500. Using the percentage method, what would be the taxes withheld for a married person who claims 3 exemptions? I wanted to redo my question

OpenStudy (anonymous):

ok so I took 35,000/52 = 673.07 then you take 70.19 x 3 = 210.57 then I did 673.07 - 210.57 = 462.50 with the table that we are using 461 - 1,455 so I took the 461 which lead me to do $462.50 – $461.00 = $1.50 then I have to take the amount of income tax to withhold is: 30.70 plus 15% $30.70 + .15 ($1.50) I came up with the finally answer as $30.70 + 0.225 = 30.72 I still believe that this is wrong some where in these steps